### v1

# Grounding Fallacies Misrepresenting Scientific Publications in Evidence

Abstract
Health-related misinformation claims often falsely cite a credible biomedical publication as evidence, which superficially appears to support the false claim. The publication does not really support the claim, but a reader could believe it thanks to the use of logical fallacies. Here, we aim to detect and to highlight such fallacies, which requires carefully assessing the exact content of the misrepresented publications. To achieve this, we introduce MissciPlus, an extension of the fallacy detection dataset Missci. MissciPlus builds on Missci by grounding the applied fallacies in real-world passages from misrepresented studies. This creates a realistic test-bed for detecting and verbalizing these fallacies under real-world input conditions, and enables novel passage-retrieval tasks. MissciPlus is the first logical fallacy dataset which pairs the real-world misrepresented evidence with incorrect claims, identical to the input to evidence-based fact-checking models. With MissciPlus, we i) benchmark retrieval models in identifying passages that support claims only when fallacies are applied, ii) evaluate how well LLMs articulate fallacious reasoning from misrepresented scientific passages, and iii) assess the effectiveness of fact-checking models in refuting claims that misrepresent biomedical research. Our findings show that current fact-checking models struggle to use relevant passages from misrepresented publications to refute misinformation. Moreover, these passages can mislead LLMs into accepting false claims as true. [^1] https://github.com/UKPLab/arxiv2024-missciplus
Grounding Fallacies Misrepresenting Scientific Publications in Evidence
Max Glockner, Yufang Hou♢♣, Preslav Nakovand Iryna Gurevych Ubiquitous Knowledge Processing Lab (UKP Lab), TU Darmstadt and Hessian Center for AI (hessian.AI) IBM Research, Ireland, MBZUAI www.ukp.tu-darmstadt.de
1 Introduction
Figure 1: We link the paraphrased context from Missci to real-world passages. The LLM must (i) find relevant passages from the original study and (ii) generate a fallacious premise to (falsely) support the claim.
Health-related misinformation has caused significant harm in our society (Zarocostas, 2020). Human fact-checking (HFC), which is time-consuming, must prioritize the most impactful claims and struggles with the rapid spread of misinformation (Arnold, 2020; Vosoughi et al., 2018). Two main approaches can automatically combat health-related misinformation. Scientific automated fact-checking (AFC) retrieves scientific evidence documents to support or to refute a claim, but faces challenges with regard to mismatched specificity (Wadden et al., 2022), and the reliance on the availability of refuting evidence (Glockner et al., 2022). Reason-checking (Musi et al., 2023) rejects claims that use fallacious reasoning, which is particularly relevant as often facts are misrepresented or skewed (Brennen et al., 2020). Most logical fallacy detection datasets require that the fallacious reasoning is explicitly stated in the text (Jin et al., 2022; Alhindi et al., 2022), and cannot detect implicit distortions from misused evidence.
The Missci dataset (Glockner et al., 2024a) addresses implicit fallacious reasoning by reconstructing the fallacious logical argument, explicitly verbalizing the flawed reasoning that led to the incorrect claim. Figure 1 shows (parts of) a fallacious logical argument from Missci for the claim that “hydroxychloroquine is a cure for COVID-19”. Each claim in Missci has a kernel of truth which is anchored in the study’s content (“Chloroquine reduced infection of the coronavirus”), denoted as the accurate premise in green. However, additional information from the study undermines the claim’s validity and reveals a reasoning gap (“The study used cell cultures for their experiments”), denoted as publication context in orange. To bridge the reasoning gap between the study content and the claim, the model must verbalize a second premise (“The results can be transferred to humans because the human body consists of cell structure.”), denoted as fallacious premise in blue. The model must also predict the type of fallacy associated with this fallacious premise.
The accurate premise and publication context in Missci constitute manually paraphrased information based on the HFC articles, which are unavailable for new claims in the wild. In a real-world scenario, models must first (i) identify the relevant passages within the entire misrepresented study and (ii) verbalize the fallacious reasoning based on the original scientific text. To address this, we present MissciPlus4), an extension of Missci with 2,257 human-annotated links between the paraphrased information from Missci and the real-world passages from the publications (see the right and the bottom of Figure 1). MissciPlus is the first fallacy dataset to pair real-world misinformation claims with publication passages as evidence. This setup, identical to the input used by AFC models, allows for the evaluation of evidence-based scientific AFC models. Since the claimant based their claim on these evidence passages, they serve as critical evidence in the fact-checking process (Silverman, 2016).
We leverage MissciPlus to answer the following research questions:
  1. 1.
    How well can existing ranking approaches select the correct evidence passages (§5)?
  2. 2.
    How well can LLMs reconstruct fallacious arguments over original studies compared to paraphrased information in Missci6)?
  3. 3.
    Can evidence-based scientific AFC models use the content of the misrepresented publications to detect the claims as misinformation (§7)?
Our findings suggest that lexical/semantic similarity based ranking models perform the best at identifying correct evidence passages. However, all models cannot leverage these passages to refute misinformation.
In summary, we contribute a new dataset MissciPlus to ground fallacies in real-world evidence, which bridges the gap to AFC. We propose novel task definitions, along with extensive experiments on retrieving the required passages of the misrepresented study, and experiments using AFC models and LLMs in detecting science distortions.
2 Related Work
Fallacy detection
research focused on surface-level fallacies (Habernal et al., 2017, 2018; Da San Martino et al., 2019; Sahai et al., 2021; Piskorski et al., 2023; Salman et al., 2023). Others extended these inventories to include logical fallacies that may require additional context for detection. All that research relied on educational examples, fake news websites (Jin et al., 2022), or fact-checking articles (Musi et al., 2022; Alhindi et al., 2022) and assumed that the explicitly stated text was sufficient to detect the fallacies. Missci (Glockner et al., 2024a) developed models to verbalize the implicit fallacious reasoning. MissciPlus differs from existing fallacy datasets as it grounds the implicit fallacies in the original content of the real-world evidence documents.
Automated scientific fact-checking
(Wadden et al., 2020; Saakyan et al., 2021; Sarrouti et al., 2021; Kotonya and Toni, 2020; Lu et al., 2023; Vladika and Matthes, 2023, 2024) used scientific documents as evidence to assess the veracity of claims. These approaches face challenges with fine-grained differences, such as specificity mismatches (Wadden et al., 2022). Our work bridges the gap between scientific AFC and fallacy detection and sheds light on the abilities of AFC models to reason over claims and distorted evidence passages.
Science communication
research (Augenstein, 2021) compares claims and the cited evidence document across various dimensions such as claim strength (Li et al., 2017), certainty of the used language (Pei and Jurgens, 2021), sentence-level causal exaggerations (Yu et al., 2020), quantification of the information match (Wright et al., 2022), or combinations of multiple dimensions (Wührl et al., 2024). Our work differs due to the focus on harmful misinformation and tasks that retrieve the passages and verbalize the fallacious reasoning.
3 Background
Preliminaries
Missci (Glockner et al., 2024a) comprises 208 health-related claims that misrepresent biomedical publications. Inspired by Cook et al. (2018), each claim is modeled as a fallacious logical argument where a claim is the conclusion of its premises. Formally, each argument comprises exactly one accurate premise ($p_{0}$) based on which the claim was made, and at least one fallacious premise $\overline{p}_{i}$ with the fallacy class $f_{i}$ (cf. §A for details of all nine fallacy classes), that is needed to (falsely) conclude the claim $\overline{c}$ from the study’s content. We refer to the pair of fallacious premise and the applied fallacy class ($\overline{p}_{i}$, $f_{i}$) as fallacy. Each fallacy in Missci is linked to one publication context ($s_{i}$) that points to a reasoning gap in the argument, and necessitates the fallacy. Each publication context (and accurate premise $p_{0}$) faithfully summarizes parts of the study. Missci defines the argument reconstruction task as: For each publication context ($s_{i}$) of the argument, given $s_{i}$, the incorrect claim ($\overline{c}$) and the accurate premise ($p_{0}$), the model must verbalize the fallacious premise and detect the applied fallacy class ($\overline{p}_{i}$, $f_{i}$). Due to overlapping fallacy definitions (Bonial et al., 2022), sometimes different interchangeable fallacies ($\overline{p}_{i}$, $f_{i}$) can bridge the same reasoning gap (expressed via $s_{i}$), and only one of them is needed to reconstruct the argument; we refer to Glockner et al. (2024a) for more detail.
In order to investigate the antiviral properties of chloroquine on SARS-CoV after the initiation of infection, Vero E6 cells were infected with the virus and fresh medium supplemented with various concentrations of chloroquine was added immediately after virus adsorption. Infected cells were incubated for an additional 16-18 h, after which the presence of virus antigens was analyzed by indirect immunofluorescence analysis. When chloroquine was added after the initiation of infection, there was a dramatic dose-dependant decrease in the number of virus antigen-positive cells (Fig. 2A). As little as 0.1-1 muM chloroquine reduced the infection by 50% and up to 90-94% inhibition was observed with 33-100 muM concentrations (Fig. 2B). At concentrations of chloroquine in excess of 1 muM, only a small number of individual cells were initially infected, and the spread of the infection to adjacent cells was all but eliminated. A half-maximal inhibitory effect was estimated to occur at 4.4 ± 1.0 muM chloroquine (Fig. 2C). These data clearly show that addition of chloroquine can effectively reduce the establishment of infection and spread of SARS-CoV if the drug is added immediately following virus adsorption. (Vincent et al., 2005).
Figure 2: A real-world passage communicates the paraphrased content $s_{1}$ from Missci that the study used cell cultures for their experiments.
Linked Passages
In Missci, the accurate premise ($p_{0}$) and publication contexts ($s_{i}$) were directly paraphrased from HFC articles and not from the original misrepresented study. Since these HFC articles were specifically written to explain to non-experts why the misrepresented study does not support the claim, the paraphrased publication contexts often reveal the reasoning gaps easily. This severely limits the applicability of Missci in the real world, where models must first identify relevant content from the misrepresented study and reason through complex scientific text. To address this, we created MissciPlus, which links the paraphrased information from Missci with actual passages from the misrepresented study. For example, Figure 2 shows a verbatim passage from the misrepresented study in MissciPlus, which is linked (i.e., communicates the same content) to the paraphrased content that “the study used cell cultures for their experiments” from Missci (see Figure 2). Formally, given a misrepresented study $S$ with its passages $S_{j}\in S$, we linked the passage $S_{j}$ to the corresponding paraphrased information ($p_{0}$ and $s_{i}$) if (parts of) $S_{j}$ entail the paraphrased information ($p_{0}$ or $s_{i}$). Any passage $S_{j}$ linked to $p_{0}$ (or $s_{i}$) can replace $p_{0}$ (or $s_{i}$) during the argument reconstruction. The same passage may be linked to multiple paraphrased information and vice versa. We denote a passage $S_{j}$ as $S_{j}^{0}$ if it links to $p_{0}$.
Subtasks
We consider three sub-tasks for reconstructing fallacious arguments in the wild. First (§5.1), the model must retrieve a passage $S_{j}^{0}$, upon which the claim is based. This is crucial for understanding the general reasoning of the claim. Second (§5.2), the model must retrieve all additional passages $S_{j}$ required to detect fallacies, i.e., passages linked to any publication context $s_{i}$. Lastly (§6), the argument reconstruction task is adapted from Missci, but replaces the paraphrased content with the respective linked passages. We always assume oracle input to establish a strong foundation for each sub-task, laying the groundwork for a more robust end-to-end system in the future.
4 Grounding Missci with Evidence
Data Collection
We selected all fallacious logical arguments from Missci, for which the misrepresented study is openly available via PMC [^2] https://www.ncbi.nlm.nih.gov/pmc/. This resulted in 118 arguments misrepresenting 100 distinct publications, which we automatically split into the constituent passages (cf. §B.1). We used the IMS model from Wright et al. (2022) to pre-select relevant passages and avoid exhaustively annotating every paraphrased information with every passage. The IMS model quantifies the information match between statements and publications, which aligns well with our needs during dataset creation. For each paraphrased information ($p_{0}$ and $s_{i}$), we selected the top-ranked passage according to IMS (cf. §B.2). We collected a minimum of six passages per argument, if possible.
Annotation
We employed two biology master’s students with annotation experience in biomedical misinformation. The annotators assigned an entailment label to all combinations of selected passage ($S_{j}$) and paraphrased information ($p_{0}$ and $s_{i}$), determining whether the passage $S_{j}$ entails (and hence is linked to) the paraphrased information. Following Glockner et al. (2024b), the annotators could express uncertainty if the entailment relation was ambiguous. If the paraphrased information was not linked to any passage, a corresponding passage was manually selected from the entire study and added to the passages for annotation. After removing four arguments for which no paraphrased information could be linked to any passage, the annotation resulted in 2,257 double-annotated entailment relations between passages and paraphrased information across 114 arguments. We retained the same instances for validation (30 arguments) and test (84 arguments) splits as in Missci. The inter-annotator agreement, measured by Cohen’s $\kappa$, was 0.602. For details about the annotation process please refer to §B.3. Overall, 400 pieces of paraphrased information (88.6% of the accurate premises $p_{0}$; 72.0% of the publication contexts $s_{i}$; 76.8% overall) could be linked to at least one corresponding passage (analysed in §B.4). We give provide descriptive statistics about the passages in §C.
5 Retrieving Relevant Passages
The passages linked to the reasoning gaps (via $s_{i}$) or to the accurate premise ($p_{0}$) are needed for the fallacious argument reconstruction and are essential for effective debunking, which must explain why a claim was thought to be true and why it actually is not (Lewandowsky et al., 2020).
5.1 Finding the Kernel of Truth
closed (annotated) open (all)
Model P@1 MRR MRR
Random 0.360 0.566 0.209
Baselines Ordered 0.480 0.658 0.443
BM25 0.547 0.705 0.539
BioBERT ST 0.547 0.712 0.582
Embeddings INSTRUCTOR 0.573 0.738 0.631
IMS 0.587 0.742 0.664
AFC (SciFact) 0.603 0.748 0.535
AFC AFC (CovidFact) 0.517 0.691 0.450
AFC (HealthVer) 0.608 0.765 0.516
AFC (all) 0.608 0.768 0.514
Llama2-70B 0.711 0.830
LLMs Llama3-8B 0.729 0.850
GPT-3.5 0.671 0.815
GPT-4 0.742 0.850
Table 1: Finding a passage $S_{j}^{0}$ based on which the claim was made (a) among the annotated passages only (closed) or (b) across the entire publication (open). Averaged over five seeds (three for LLMs).
Given an incorrect claim $\overline{c}$ that misrepresents a publication $S=[S_{0},S_{1},...,S_{n}]$, the model must rank all passages $S_{i}\in S$ such that the top-ranked passage communicates the information ($p_{0}$) based on which the false claim $\overline{c}$ was made (denoted as $S_{j}^{0}$). This task is similar to finding supporting evidence passages in automated fact-checking (Thorne et al., 2018), but it differs in that the evidence passage is in a “corrupted” support relationship with the claim, meaning it only supports the claim when a fallacy is involved. Finding the passage $S_{j}^{0}$, which explains the basis of the claim, is crucial as it reveals the (broken) rationale behind the claim. We report P@$1$ and MRR over the subset of annotated passages with comprehensive annotations (closed) and over all passages (open). The open evaluation is more realistic, but only serves as a lower bound.
5.1.1 Models
As baselines, we randomly shuffled passages (random) or maintained their original order from the publication (ordered). We used BM25 for lexical similarity-based ranking. For semantic similarity, we ranked the passages by their cosine similarity to the claim using sentence embeddings from BioBERT (Lee et al., 2020), fine-tuned by Deka et al. (2022) for evidence selection in scientific AFC, and prompt-based embeddings from INSTRUCTOR (Su et al., 2022). We also report the performance of the IMS (Wright et al., 2022) used during the dataset construction. For cross-encoders, we trained DeBERTaV3 (He et al., 2022) models on three scientific AFC datasets SciFact (Wadden et al., 2020), CovidFact (Saakyan et al., 2021), HealthVer (Sarrouti et al., 2021), and their union (all) (cf. §D). Given the claim and the passage, we compute and rank passages based on the predicted label probability for Supported. This label fits closest to the task definition of finding a passage that seemingly supports the claim. We used LLMs only in the closed evaluation due to their computational cost. Here, we implemented PRP (Qin et al., 2023), which reorders passages through pairwise comparisons akin to the early iterations of the bubble sort algorithm. We used GPT-4 (OpenAI, 2023) and GPT-3.5333https://platform.openai.com/docs/models/ as commercial LLMs, Llama3-8B [^4] https://llama.meta.com/llama3/ and Llama2-70B (Touvron et al., 2023) as open-source LLMs. The implementation details are outlined in §E.
5.1.2 Results
Table 1 shows solid performance across all evaluated models. The strong ordered baseline suggests that claims often rely on early parts of a study. In the open evaluation, embedding-based approaches perform the best. Note that the IMS model preselected the passages for annotation (in a simplified setup that provided the accurate premise instead of the claim), and its performance must be interpreted with caution. AFC models are superior to embedding models in the closed evaluation, but fall behind PRP ranking via LLMs.
5.2 Finding Undermining Passages
Given an incorrect claim $\overline{c}$ that misrepresents a publication $S=[S_{0},S_{1},...,S_{n}]$, and a passage $S_{j}^{0}\in S$, based on which the claim was made, the model must rank all passages $S_{i}\in S$ such that the top-ranked passages $S_{i}$ are linked to reasoning gaps (expressed via the publication context $s_{i}$) between $S$ and $\overline{c}$. This passage ranking task starkly differs from evidence retrieval in AFC because the model must reason about the relationships between the claim, the accurate premise ($p_{0}$) on which the claim is based, and the content of the passage. For example, in vitro experiments with cell cultures only indicate a fallacy because the claim relies on the results of these experiments ($p_{0}$) but incorrectly suggests that they indicate a cure for humans. Additionally, this task seeks content that reveals a fallacy by disrupting the reasoning, which is different from retrieval for multi-hop reasoning that connects multiple pieces of information to establish a reasoning path. Since multiple passages are often needed to find all fallacies, we report the passage-level mean average precision (MAP) as our main metric. We further report P@$1$ because one detectable fallacy is the minimum requirement to reject a claim. To measure how many distinct reasoning gaps can be detected, we report the fallacy-level recall of the ten top-ranked passages (Fall-R@$10$). This penalizes models that only detect passages linked to the same $s_{i}$. We only evaluate on the open subset, as most annotated passages are linked with reasoning gaps as per the dataset construction.
Figure 3: Recall of undermining passages per fallacy class (and accurate premise) over the top $k$ ranked passages. We only list fallacies with $\geq 20$ occurrences.
5.2.1 Models
We use the ranking models from §5.1. We slightly adapted the AFC-based ranking to use the sum of the predicted probabilities for the labels Supported and Refuted. We further assume oracle outputs from the previous subtask and prepended a randomly sampled gold $S_{j}^{0}$ to the claim in all baselines. These design choices follow our experiments on the validation split (cf. §F).
5.2.2 Results
Model MAP P@1 Fall-R@10
Random 0.205 0.136 0.438
Baselines Ordered 0.286 0.298 0.497
BM25 0.496 0.617 0.602
INSTRUCTOR 0.541 0.652 0.613
Embedding SPICED-IMS 0.524 0.640 0.595
BioBERT ST 0.491 0.600 0.570
AFC (SciFact) 0.360 0.326 0.518
AFC AFC (CovidFact) 0.380 0.457 0.538
AFC (HealthVer) 0.368 0.410 0.544
AFC (all) 0.306 0.338 0.554
Table 2: Retrieving passages linked to fallacies from the entire publication (open). Averaged over five seeds.
To solve this task, the model must first understand the (false) rationale behind the claim as expressed in the passage $S_{j}^{0}$ (e.g., communicating that “chloroquine reduced infection of the coronavirus”). Then, the model can identify how a different passage interacts with this reasoning to highlight a reasoning gap (e.g., that this was observed in “in vitro” experiments). Table 2 shows that cosine-similarity-based ranking outperforms all AFC-based models that can jointly reason over the claim and the evidence despite the complex reasoning that is required. This suggests that lexical or semantic similarities correlate sufficiently strongly with passages that indicate reasoning gaps, while AFC models seemingly fail to reason over them.
To understand whether fallacy classes impact the difficulty of passage retrieval, we show the recall of detectable fallacy classes over the number of retrieved passages for the INSTRUCTOR and AFC (all) in Figure 3. Overall, within the same model, different fallacy classes follow similar trends. The superior performance of INSTRUCTOR seemingly relies on fallacies that can be detected from $S_{j}^{0}$ passages. For comparison, we visualize the performance without $S_{j}^{0}$ passages in §F.2, Figure 9, which reduces the performance gap between the two ranking models.
6 Fallacious Argument Reconstruction
Given an incorrect claim $\overline{c}$ and relevant selected passages $S_{j}\in S$ of the misrepresented study, the model must verbalize the fallacious premises $\hat{\overline{p}}_{i}$ with the fallacy class $\hat{f}_{i}$ such that all reasoning gaps between $S$ and $\overline{c}$ can be bridged via the generated fallacious premises $\hat{\overline{P}}=[\hat{\overline{p}}_{1},\hat{\overline{p}}_{2},...,\hat{\overline{p}}_{n}]$. We assume oracle output of both passage retrieval tasks and only consider passages linked to a fallacy, ensuring that at least one passage is linked to the accurate premise (as the basis for the claim). We experimented with GPT-4 Turbo as a proprietary LLM, GPT-3.5 as a lower-cost alternative, and Llama3-8B as a representative small open-source LLM with strong leaderboard performance. As in Missci, we prompted the LLMs in a zero-shot setting (cf. §G.1 for prompt selection). Figure 4 shows examples for the Definition, Logical Form and toy Examples for the Fallacy of Composition taken from literature (Bennett, 2012; Cook et al., 2018). By default, all prompts include the these properties as supplementary information for each fallacy (cf. §A for all fallacy classes).
Fallacy class: Fallacy of Composition
Definition: Inferring that something is true of the whole from the fact that it is true of some part of the whole
Logical Form: A is part of B. A has property X. Therefore, B has property X.
Example: Hydrogen is not wet. Oxygen is not wet. Therefore, water (H2O) is not wet.
Figure 4: Examples for the (D)efinition, (L)ogical form and (E)xample for the Fallacy of Composition, used as supplementary fallacy information in the prompts.
Missci MissciPlus
LLM Info R@5 ($\phi^{\mathrm{f+p}}$) R@5 ($\phi^{\mathrm{f}}$) Arg@$1$ ($\phi^{\mathrm{f+p}}$) R@5 ($\phi^{\mathrm{f+p}}$) R@5 ($\phi^{\mathrm{f}}$) Arg@$1$ ($\phi^{\mathrm{f+p}}$)
DLE 0.277 0.514 0.552 0.226 0.477 0.476
Llama3-8B DL 0.241 0.445 0.512 0.195 0.463 0.413
DE 0.227 0.470 0.480 0.174 0.449 0.389
LE 0.255 0.469 0.504 0.209 0.439 0.460
DLE 0.248 0.491 0.512 0.165 0.428 0.361
GPT-3.5 DL 0.232 0.492 0.464 0.146 0.416 0.321
DE 0.276 0.517 0.567 0.160 0.400 0.333
LE 0.249 0.478 0.524 0.157 0.410 0.341
DLE 0.332 0.486 0.619 0.224 0.458 0.452
GPT-4 Turbo DL 0.308 0.500 0.583 0.238 0.495 0.488
DE 0.318 0.528 0.595 0.210 0.491 0.440
LE 0.304 0.505 0.583 0.252 0.519 0.500
Table 3: Argument reconstruction performance using paraphrased information from Missci compared to the real passages from MissciPlus across various fallacy information. Results (expect GPT-4) are averaged over 3 seeds. We evaluate different combinations of fallacy (D)efinition, (L)ogical form and (E)xample in the prompt.
LLM Passages R@5 P@1 Arg@1
Llama3-8B per-passage 0.226 0.290 0.476
Llama3-8B all passages 0.199 0.266 0.425
GPT-3.5 per-passage 0.165 0.190 0.361
GPT-3.5 all passages 0.089 0.194 0.206
Table 4: All passage prompting vs. per passage prompting. Averaged over three seeds.
6.1 Holistic Argument Evaluation
The evaluation on MissciPlus faces two challenges: First, in Missci, each reasoning gap was expressed by one publication context ($s_{i}$) for which dedicated fallacious premises exist. LLMs were prompted for each $s_{i}$ separately, and the predicted fallacious could be directly compared to the annotated fallacies for the same $s_{i}$. MissciPlus does not have a direct mapping between the evidence passage $S_{j}$ and the dedicated fallacies since the same passage can be linked to multiple reasoning gaps ($s_{i}$) and vice versa. Second, the arguments can contain different reasoning gaps ($s_{i}$) that can be bridged with different fallacious premises that apply the same fallacy class. For example, the claim that “hydroxychloroquine is a cure for COVID-19” in Figure 1 falsely assumes (among others) that i) hydroxychloroquine will have an equal effect as chloroquine, and that ii) SARS-CoV-2 will behave equally to SARS-CoV-1, both of which use the False Equivalence fallacy, but address distinct gaps. This further complicates matching the generated fallacies to the predicted ones. To address this, we evaluated the fallacious argument holistically: Given all relevant passages of the misrepresented study, we expect up to five ranked fallacies ($\hat{\overline{p}}_{i}$, $\hat{f}_{i}$), where five is the maximum number of distinct reasoning gaps in Missci. Following (Schlichtkrull et al., 2023), we automatically match the generated fallacies to the gold fallacies at the argument level (instead of using $s_{i}$ as in Missci).
Evaluation and Matching
Given a predicted fallacy $\hat{y}_{i}$ = ($\hat{\overline{p}}_{i}$, $\hat{f}_{i}$) and a gold fallacy $y_{k}$ = ($\overline{p}_{k}$, $f_{k}$), we define a function $\phi:(y,\hat{y})\rightarrow\{0,1\}$ which discerns if the predicted and the gold fallacy are a match based on two implementations:
  • $\phi^{\mathrm{f}}$ outputs $1$ if the predicted fallacy class equals the gold fallacy class.
  • $\phi^{\mathrm{f+p}}$ outputs $1$ if the generated fallacious premise additionally bridges the same reasoning gap as the gold fallacious premise.
The $\phi^{\mathrm{f}}$ is an upper bound, as it does not penalize models for poorly phrased premises. $\phi^{\mathrm{f+p}}$ uses a Llama3-8B model fine-tuned with QLoRA adapters (Dettmers et al., 2023) on the human evaluation data from Missci (cf. §G.2). The performance via cross-validation is 79.8 points in accuracy (78.8 in F1-macro), suggesting that it is suitable, albeit imperfect, in matching the fallacious premises. We do not perform a human evaluation, which is too complex given the many-to-many relationship between the predicted and the gold fallacies, and is hard to reproduce for future work. Instead, we report evaluation measures based on the two complementary implementations $\phi$, which we deem adequate to answer our research questions. As a primary measure, we report the recall of reasoning gaps for which $\phi$ found a match among the five fallacy predictions (R@$5$). Following Glockner et al. (2024a), we use precision P@$1$, which checks whether the top-ranked fallacy is correct based on $\phi$, and Arg@$1$, which considers an argument as successfully rejected if at least one of the predicted fallacies is correct based on $\phi$.
6.2 Experiments
6.2.1 Benefit of Per-Passage prompting
We compare the performance when prompting LLMs based on each passage $S_{j}$ individually (per-passage) and when including all concatenated passages in a single prompt (all passages). The per-passage prompts follow Missci, where the fallacies where detected for each publication context ($s_{i}$) with the provided accurate premise ($p_{0}$). To apply this on MissciPlus, we replaced the paraphrased publication context $s_{i}$ with a passage $S_{j}$ that is linked to $s_{i}$, and replace $p_{0}$ with a randomly sampled passage $S_{j}^{0}$. For both prompting strategies, we selected the five top-ranked fallacy predictions per argument for evaluation using $\phi^{\mathrm{f+p}}$ in Table 4.
Per-passage prompting is superior for both LLMs, but more expensive due to multiple prompts per argument. The higher (or similar) precision suggests that focusing on one passage (per-passage prompting) can be advantageous when identifying the fallacies. Llama3-8B outperforms GPT-3.5 across all measures, but detects none of the annotated fallacies for more than half of the misinformation, according to the automated metrics.
6.2.2 Missci vs. MissciPlus
Table 3 compares the performance of all three LLMs and ablations over various fallacy information in the prompts (Definition, Logical Form and Explanation) on Missci and MissciPlus using per-passage (per-publication context $s_{i}$ on Missci) prompting. We can see that GPT-4, as a state-of-the-art LLM, performs best across all measures and datasets on this task. We further observe a considerable impact of different fallacy information for each LLM. Yet, none is universally beneficial or harmful for all tested LLMs.
A clear trend of decreasing performance from the paraphrased information in Missci to the real-world passages in MissciPlus is evident. We note that, the approximate measures via $\phi^{\mathrm{f}}$ and $\phi^{\mathrm{f+p}}$ may underestimate performance as they cannot match valid fallacies not covered by the annotations (Glockner et al., 2024a). The performance from Missci to MissciPlus only drops marginally for Llama3-8B and GPT-4 Turbo. Interestingly, GPT-3.5 outperforms Llama3-8B using the paraphrased content. However, its poor performance across all evaluated prompts on MissciPlus corroborates the findings in Table 4 and suggests poor adaptability toward realistic scientific text on this task.
7 Scientific AFC Evaluation
AFC models rely on retrieved evidence to refute misinformation (Guo et al., 2022). However, evidence that directly refutes a claim may not exist (Glockner et al., 2022), and misinformation often rather distorts the facts (Brennen et al., 2020). MissciPlus contains paired claims with real-world evidence passages that can be assessed by fallacy detection models and evidence-based fact-checking models. Using MissciPlus, we now test how well AFC models and LLMs can detect the nuanced real-world misrepresentation of scientific biomedical publications.
7.1 Fine-Tuned Models for Scientific AFC
Argument level AFC
Training Data Sup. Ref. Mix. NEI Acc.
SciFact 55.5 4.5 23.3 16.7 88.9
HealthVer 39.0 20.0 27.1 13.8 82.1
CovidFact 36.9 12.1 51.0 90.7
All AFC 48.8 11.9 27.1 12.1 84.1
Table 5: Veracity predictions from scientific AFC models on 84 misinformation claims with 510 evidence passages in MissciPlus. Averaged over five seeds.
Given a claim $\overline{c}$ that misrepresents the publication $S$, we form $n$ fact-checking instances ($\overline{c}$, $S_{j}$) by treating each of the $n$ annotated passages $S_{j}\in S$, as provided evidence. We use the scientific AFC models used from §5 to produce $n$ veracity labels given $\overline{c}$ and each of the selected passages $S_{j}$. Following Schlichtkrull et al. (2023), we assign an overall veracity label for a claim as Mixed if the labels Supported and Refuted were predicted. If the model only predicted NotEnoughInformation (NEI), the overall verdict was NEI. In all other cases in which the model predicted a mix of Supported (with optional NEI) or Refuted (with optional NEI) we label it as Supported or Refuted respectively. We note that the studies in MissciPlus constitute trustworthy evidence. They do not support the inaccurate claims that HFC has rated as misrepresenting these studies. A literate scientific AFC model should equally detect the claims as misinformation based on the same trustworthy evidence.
Figure 5: AFC predictions over passages linked to the accurate premise (top left), to reasoning gaps (top right), to both (bottom left) or none (bottom right).
Table 5 shows that all AFC models yielded high in-domain accuracy on the respective AFC dataset (AFC Acc.) ranging from 82-90 points in accuracy. Because Missci assumes that the study provides sufficient evidence for the claim, any prediction other than Supported can be considered as correctly rejecting the claim. Yet, AFC systems falsely predict 37-56% of the claims to be true. The seemingly best performing CovidFact model uses binary classification without NEI and benefits from a reduced chance to predict the overall label Supported as per the overall label assignment.
The evaluation on misinformation favors sceptical models regardless of whether a passage truly undermined a claim. Grounding the verdict of AFC models is critical for the trustworthiness and often part of the evaluation protocol (Thorne et al., 2018; Wadden et al., 2020; Glockner et al., 2024b). To understand whether the AFC models reject the claims based on the undermining evidence that indicates reasoning gaps, we visualize the AFC prediction over different passages in Figure 5. Intuitively, the models mostly predict Supported over passages based on which the claim was made (top left). If the same passage is additionally linked to a reasoning gap (bottom left), the predicted distribution does not change much, suggesting unawareness of the mismatches. Based on passages that are linked to a reasoning gap but not to the accurate premise (top right), the distribution changes towards NEI. This is similar to completely unrelated passages (bottom right) and suggests that AFC models struggle to detect scientific distortions as such. We provide additional analysis in §H.1.
Predicted as
LLM True False NEI True False NEI
Llama2 1.6 61.1 37.3 34.7 22.3 41.3
Know Llama3 0.0 86.9 2.4 20.0 43.3 14.3
(FC) GPT4 0.0 85.3 14.7 59.0 23.0 17.3
GPT3.5 1.2 71.0 28.2 46.7 17.3 35.7
Llama2 0.0 100.0 0.0 29.7 69.3 3.0
Know Llama3 8.3 88.9 2.4 68.7 26.0 3.7
(Ask) GPT4 3.6 68.3 27.8 49.7 36.3 6.7
GPT3.5 1.6 50.4 48.0 47.3 6.0 45.0
Llama2 23.8 61.5 12.7 58.7 29.7 10.7
RAG Llama3 44.4 53.2 2.4 80.3 16.3 3.3
GPT4 27.4 34.1 38.5 55.0 4.0 41.0
GPT3.5 38.9 31.7 29.0 78.0 5.3 16.0
Misinformation True claims
Table 6: Averaged veracity predictions from LLMs on misinformation from MissciPlus (left) and true claims from HealthVer and CovidFact (right).
7.2 LLMs as Scientific AFC Models
Experimental Setup
We tested how well LLMs can fact-check claims in MissciPlus when prompted to predict their veracity directly. We prompted each LLM (a) without any evidence to test whether it already has parametric knowledge about the claims from pretraining (Magar and Schwartz, 2022) and (b) with the relevant evidence passages to test their reasoning capabilities in a RAG style, similarly to Manakul et al. (2023). For parametric knowledge, we let the LLM generate a fact-checking article (FC), which is closest to our application, but may bias the model due to the prevalence of misinformation in fact-checking articles. As a neutral prompt, we additionally asked the LLM to predict the veracity to the best of their knowledge. All prompts are listed in §I.
Results
Table 6 shows the results on MissciPlus and 100 randomly sampled correct claims from HealthVer and CovidFact (50 from each dataset), both of which focus on real-world health-related claims from the internet. Note that the percentages do not always add up to 100% as LLMs can refuse to answer. The results suggest that all LLMs seem to have learned that the claims in MissciPlus are not true. Also, true claims are rarely classified as False and mostly as True. The proportion of claims predicted as True increases when providing additional evidence. Interestingly, this also happens to misinformation in MissciPlus despite the LLM seemingly knowing that these claims are not true and despite that the misrepresented evidence does not truly support the claims. This is similar to overruling the internal knowledge with evidence (Wu et al., 2024) and suggests that the LLMs do not identify the reasoning gaps between the claim and the evidence, as outlined by the HFCs. Our findings indicate that LLMs, similarly to humans, are prone to misinterpreting scientific publications. Due to their high persuasiveness (Augenstein et al., 2023; El-Sayed et al., 2024), this can lead to disastrous problems when people rely on LLMs, even if LLMs transparently output the (misrepresented) used evidence.
Discussion
Similarly to previous work on fallacy recognition (Jin et al., 2022; Alhindi et al., 2022; Glockner et al., 2024a), the prompts in this work focus on which fallacies apply instead of whether any applies, and are no panacea either: In preliminary experiments §H.2, we prompt LLMs to verbalize the fallacies or output that there is “no fallacy”. The LLMs found fallacies in 85-99% of the claims in MissciPlus but also in 78-99% of the true claims. MissciPlus unifies fallacy detection with AFC and provides a resource for studying how to handle fallacious arguments and true claims alike.
8 Conclusion and Future Work
We introduced MissciPlus, a new dataset for reconstructing fallacious arguments based on realistic evidence. We showed that existing ranking models and LLMs struggle to reconstruct fallacious arguments using the evidence that reveals reasoning gaps. Moreover, fine-tuned AFC models and LLMs failed to refute claims in MissciPlus when presented with misrepresented evidence, highlighting the dangers of persuasive LLMs.
In future work we plan to leverage MissciPlus to build and evaluate models that are resistant to the nuanced mismatches between claims and their corresponding cited evidence documents.
Limitations
MissciPlus assumes that the claims are based on a single publication only and that each publication is inherently trustworthy, i.e., that the only error was done in the reasoning from the publication to the claim. In the real world, finding complementary credible evidence is critical. We found that single passages are insufficient to ground each fallacy in evidence and ignore content that is multi-model, requires multiple passages or can only be found from the entire study. Our evaluation is based on automatic matching, which is inevitably imperfect. Models may detect fallacies that are valid but not contained in the original annotations in Missci. We compensate for this with two complementary matching strategies and experiments over different seeds and prompt variations to confirm the robustness of our observations. By focusing on recall, we further do not penalize models for prediction fallacies that are not within our annotation while still requiring the models to detect the most prominent ones as highlighted by the annotators in Missci (from the HFC articles).
Ethics Statement
The research questions targeted in this work aim to improve the detection of claims that distort scientific publications, which is ethically uncritical. Ethical concerns are bound to cases in which the content of this study are used in unintended ways.
Dual Use
False interpretations of health-related claims can have disastrous effect. Any output of models derived from MissciPlus only serves research purposes to detect such misinformation, but under no circumstances must be considered accurate without consulting experts in the field. Our work poses dangers for dual use, particularly verbalizing the fallacious reasoning to draw incorrect conclusions from real-world studies. While generating (parts of) misinformation always poses a risk, it is unavoidable to build resilience against real-world misinformation, as demonstrated in previous work (Zellers et al., 2019; Huang et al., 2023; Alhindi et al., 2023; Glockner et al., 2024a).
Data Collection
All publications used in MissciPlus have been published by the respective authors and we did not anonymize their work. All publications used in MissciPlus are openly available and are part of the public discourse; in fact, they have even been distorted by misinformation. Hence, similarly to other scientific corpora (Lehman et al., 2019; Lo et al., 2020) that rely on such publications as evidence, we did not ask for explicit permission from the authors of each study to use their work..
Acknowledgments
This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE and by the German Research Foundation (DFG) as part of the UKP-SQuARE project (grant GU 798/29-1). Yufang Hou is supported by the Visiting Female Professor Programme from TU Darmstadt. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). We are grateful to our dedicated annotators who helped to create MissciPlus. Finally, we wish to thank Jan Buchmann, Sukannya Purkayastha, Luke Bates and Jonathan Tonglet for their valuable feedback on an early draft of this work.
Appendix A Fallacies in MissciPlus
Definition Logical Form
Ambiguity
When an unclear phrase with multiple definitions is used within the argument; therefore, does not support the conclusion. Claim X is made. Y is concluded based on an ambiguous understanding of X.
Equivocation (merged with Ambiguity)
When the same word (here used also for phrase) is used with two different meanings. Equivocation is a subset of the ambiguity fallacy. Term X is used to mean Y in the premise. Term X is used to mean Z in the conclusion.
Impossible Expectations / Nirvana Fallacy
Comparing a realistic solution with an idealized one, and discounting or even dismissing the realistic solution as a result of comparing to a “perfect world” or impossible standard, ignoring the fact that improvements are often good enough reason. X is what we have. Y is the perfect situation. Therefore, X is not good enough.
False Equivalence
Assumes that two subjects that share a single trait are equivalent. X and Y both share characteristic A. Therefore, X and Y are [behave] equal.
False Dilemma
Presents only two alternatives, while there may be another alternative, another way of framing the situation, or both options may be simultaneously viable. Either X or Y is true.
Biased Sample Fallacy
Drawing a conclusion about a population based on a sample that is biased, or chosen in order to make it appear the population on average is different than it actually is. Sample S, which is biased, is taken from population P. Conclusion C is drawn about population P based on S.
Hasty Generalization
Drawing a conclusion based on a small sample size, rather than looking at statistics that are much more in line with the typical or average situation. Sample S is taken from population P. Sample S is a very small part of population P. Conclusion C is drawn from sample S and applied to population P.
False Cause Fallacy (use as Causal Simplification)
Post hoc ergo propter hoc — after this therefore because of this. Automatically attributes causality to a sequence or conjunction of events. A is regularly associated with B; therefore, A causes B.
Single Cause Fallacy (use as Causal Simplification)
Assumes there is a single, simple cause of an outcome. X is a contributing factor to Y. X and Y are present. Therefore, to remove Y, remove X.
Fallacy of Composition
Inferring that something is true of the whole from the fact that it is true of some part of the whole. A is part of B. A has property X. Therefore, B has property X.
Fallacy of Division (merged with Fallacy of Composition)
Inferring that something is true of one or more of the parts from the fact that it is true of the whole. A is part of B. B has property X. Therefore, A has property X.
Fallacy of Exclusion / Cherry Picking / Slothful Induction
When only select evidence is presented in order to persuade the audience to accept a position, and evidence that would go against the position is withheld (Cherry Picking). Ignores relevant and significant evidence when inferring to a conclusion (Slothful Induction – focus on neglect). Evidence A and evidence B is available. Evidence A supports the claim of person 1. Evidence B supports the counterclaim of person 2. Therefore, person 1 presents only evidence A.
Table 7: Fallacy Overview. Definition and logical form taken from Bennett (2012) and Cook et al. (2018). Table as provided by Glockner et al. (2024a).
Ambiguity
It is said that we have a good understanding of our universe. Therefore, we know exactly how it began and exactly when.
Equivocation
A feather is light. What is light cannot be dark. Therefore, a feather cannot be dark.
Impossible Expectations / Nirvana Fallacy
Seat belts are a bad idea. People are still going to die in car crashes.
False Equivalence
They are both Felidae, mammals in the order Carnivora, therefore there’s little difference between having a pet cat and a pet jaguar.
False Dilemma
I thought you were a good person, but you weren’t at church today.
Biased Sample Fallacy
Based on a survey of 1000 American homeowners, 99% of those surveyed have two or more automobiles worth on average \$100,000 each. Therefore, Americans are very wealthy.
Hasty Generalization
My father smoked four packs of cigarettes a day since age fourteen and lived until age sixty-nine. Therefore, smoking really can’t be that bad for you.
False Cause Fallacy
Every time I go to sleep, the sun goes down. Therefore, my going to sleep causes the sun to set.
Single Cause Fallacy
Smoking has been empirically proven to cause lung cancer. Therefore, if we eradicate smoking, we will eradicate lung cancer.
Fallacy of Composition
Hydrogen is not wet. Oxygen is not wet. Therefore, water (H2O) is not wet.
Fallacy of Division
His house is about half the size of most houses in the neighborhood. Therefore, his doors must all be about 3 1/2 feet high.
Fallacy of Exclusion / Cherry Picking / Slothful Induction
Employer: “It says here on your resume that you are a hard worker, you pay attention to detail, and you don’t mind working long hours.”
Andy: “Yes sir.”
Employer: “I spoke to your previous employer. He says that you constantly change things that should not be changed, you could care less about other people’s privacy, and you had the lowest score in customer relations.”
Andy: “Yes, that is all true, as well.”
Table 8: Fallacy Examples (taken from Bennett (2012)). Table as provided by Glockner et al. (2024a).
We provide the definitions, logical form, and examples from literature taken from Glockner et al. (2024a) in Tables 7-8. Missci aggregates the fallacies into nine different classes:
  1. 1.
    Ambiguity: Combines Ambiguity and Equivocation.
  2. 2.
    Impossible Expectations: Uses Impossible Expectations and its alternative names.
  3. 3.
    False Equivalence: Uses False Equivalence.
  4. 4.
    False Dilemma: Uses False Dilemma.
  5. 5.
    Biased Sample Fallacy: Uses Biased Sample Fallacy.
  6. 6.
    Hasty Generalization: Uses Hasty Generalization.
  7. 7.
    Causal Simplification: Combines False Cause Fallacy and Single Cause Fallacy.
  8. 8.
    Fallacy of Composition: Combines Fallacy of Composition and Fallacy of Division.
  9. 9.
    Fallacy of Exclusion: Uses Fallacy of Exclusion and its alternative names.
Appendix B Dataset Construction
B.1 Passage Extraction
When documents are available in full-text via the PMC API [^5] See https://www.ncbi.nlm.nih.gov/books/NBK25499/#chapter4.EFetch, we retrieve the XML document and collect text passages enclosed within <p> tags within the <body> node. If the full-text document is accessible as HTML but not through the API, we extract the HTML content from the Wayback Machine [^6] https://archive.org/web/ and separate the passages based on HTML structure. We make the data collection script publicly available.
B.2 Passage Selection
We segment each passage into constituent sentences using SciSpacy (Neumann et al., 2019) and compute the passage-level IMS (Wright et al., 2022) by max-pooling the IMS of all sentences. We retain the passage with the highest max-pooled IMS for each argument paraphrased information that needs a match ($p_{0}$, $s_{i}$). More passages (if available) are selected based on the highest IMS until we have six passages per argument. The IMS is particularly well-suited for our problem as it quantifies the information of the scientific findings rather than focusing on semantic similarity
B.3 Annotation
Figure 6: Annotation Interface from Surge AI to assign the entailment labels between the a passage and the paraphrased information.
Stage Args Passages Relations Mapped
1st Round 118 681 2,191 64.3%
2nd Round 118 719 2,334 67.9%
Consolid. 118 719 2,334 74.2%
Cleaned 114 694 2,257 76.8%
Table 9: Stages of the argument mapping annotation, including how many of the paraphrased information are linked to passages.
We use Surge AI [^7] https://www.surgehq.ai/ as the annotation platform and employed two M.Sc. students in biology paid 12.26 EUR per hour. We provide a screenshot of the annotation interface in Figure 6 and construct MissciPlus in multiple rounds. Table 9 provides statistics for each round of dataset construction.
Passage Linking
First, every combination of paraphrased information ($s_{i}$ or $p_{0}$) and selected passage $S_{j}$ is double annotated by our annotators, deciding whether (parts of) the passage entail the paraphrased information. To account for cases without a clear entailment label, we follow Glockner et al. (2024b) and allow annotators to express their uncertainty. Specifically, annotators can choose one of the labels “entailed”, “probably entailed”, “probably not entailed” or “not entailed”. We conservatively only consider the link between paraphrased information and passaged and as entailed if at least one of both annotators labeled it as “entailed” and the other labeled it as either “entailed” or “probably entailed”. In the first round, we annotated 2,191 relations with 681 distinct passages. This annotation round linked 266 paraphrased content to at least one passage. In a second round, one annotator manually identified (if possible) a corresponding passage from all remaining passages for each missing link. These new passages were then double annotated with every paraphrased information ($s_{i}$, $p_{0}$) as outlined before. This annotation round contained 38 additional passages with 143 new relations and increased the number of paraphrased information linked to a passage to 281. The agreement for the fine-grained labels is 0.445. When merging “entailed” with “probably entailed” and “not entailed” with “probably not entailed”, the inter-annotator agreement rises to 0.602. This suggests that some of the disagreement stems from uncertainty regarding a definite label.
Consolidation
To account for possible false negatives due to our conservative label aggregation, we consolidate annotations where the two annotators reached no unanimous label. We consider instances as not unanimous if they include instances labeled as “probably entailed” by both annotators or “entailed” by one and “probably not entailed” or “not entailed” by the other. To finalize the overall label, we provided the annotator with the entire fallacious argument during consolidation. This additional context allowed for a more contextual understanding of the role of the passage in reconstructing the fallacious argument. Our primary annotator, who had extensive experience in fallacy annotation and was involved in all our pilot studies, handled consolidation. Cases in which our consolidator could not clearly identify as “entailed” were labeled as “not entailed”. In total, 245 relations needed consolidation, 58 of which were consolidated as “entailed”, leading to 26 previously unlinked paraphrased information. Four arguments were not linked to any passages and were subsequently removed. This likely happened when the link to the misrepresented publication was falsely selected in Missci.
B.4 Missing Passage Link Analysis
# Sentence
(1) The study did not include a group that did not wear masks at all. (negated)
(2) The experiments were done on concentrations that are different from concentrations found in patients or vaccinated people. (scope)
(3) Chloroquine diphosphate and hydroxychloroquine sulfate show antiviral activity against MERS-CoV and SARS-CoV. (multi-hop)
(4) The average amount of spike protein in the blood was about 30 to 40 picograms/mL after receiving the Moderna vaccine. (multi-modal)
Table 10: Examples of paraphrased information with no linked passage.
During dataset construction, we assumed that each paraphrased information from Missci could be linked to a single passage. This was feasible in 400 cases only (76.8%). Accurate premises were more frequently linked to at least one passage (88.6%), than publication contexts (72.0%). We list representative reasons where no passages could not be found in Table 10. The most common reason, accounting for 41.6%, was the presence of negation in the paraphrased information, discussing information not present in the study. For instance, a claim questioning mask effectiveness would require a study with a control group without masks. However, if the study did not focus on mask effectiveness in general, there is no need for such a control group or to explicitly mention its absence. Negated sentences were more prevalent in undermining publication contexts (47.4%) and less frequent in accurate premises (7.7%). Among the remaining instances, we identified (2) scope mismatches between the claim and the study, (3) information spread across multiple passages, and (4) non-textual components like tables or figures. While these fields could not be linked to a single passage, it is theoretically possible to reconstruct the argument using the complete publication.
B.5 Licenses
Missci extends the Missci dataset, published under the Apache 2.0 license. MissciPlus aligns directly with their intended use, which is to outline the fallacious reasoning of distorted science transparently, and only improved its applicability in the real world. Our collected annotations and all scripts to collect and preprocess the scientific publications will be made publicly available under an open-source license. All fallacious logical argument and all publications in MissciPlus are in English.
Appendix C Length of Passages in MissciPlus
Figure 7: Number of sentences per passage over all annotated passages in MissciPlus. Passages are only considered once (if used by multiple arguments).
Figure 8: Number of sentences per passage over all passages of the entire publication in MissciPlus. Passages are only considered once (if used by multiple arguments).
The number of annotated passages per argument ranges from 1-8 passages (mean: 6.1; std: 1.2; median: 6.0). The number of all annotations per argument ranges from 1-114 passages (mean: 43.7; std: 25.5; median: 43.0). Annotated passages vary in length, with 1 to 28 sentences. These passages contain 4.9 sentences on average (median: 4). When considering all passages of the entire misrepresented publication, each passage averages 3.8 sentences, with a median of 3. Figure 7 displays the distribution of sentence counts in annotated passages, while Figure 8 shows the same for all passages.
Appendix D Scientific AFC Models
We train four DeBERTaV3 (He et al., 2022) models for scientific AFC using SciFact (Wadden et al., 2020), CovidFact (Saakyan et al., 2021), HealthVer (Sarrouti et al., 2021) and the combination of all.
D.1 Scientific AFC datasets
SciFact (Wadden et al., 2020) covers multiple scientific domains using scientific abstracts as evidence documents. Citations were used to collect claims supported by the evidence. Refuted claims were generated via careful paraphrasing. While the original task includes the selection of evidence documents and rationale sentences, we only train models to predict the stance towards the claim given the full abstract as evidence. We used the official validation split as our test split (due to the hidden official test split). We randomly selected 200 instances from the official training set as the validation set and trained on the remaining instances. CovidFact (Saakyan et al., 2021) covers scientific claims related to COVID-19. Pairs of claims and evidence originate from a strictly moderated Reddit forum where every claim must provide an evidence document. Refuted claims were created by automatically changing the true claims. Unlike other AFC datasets, this dataset only distinguishes between “supported” and “refuted” claims, missing a label for “not enough information”. HealthVer (Sarrouti et al., 2021) used COVID-19-related questions as queries and collected claims from resulting web pages via a search engine. Each claim was annotated against retrieved scientific abstracts as evidence using the three labels Supports, Refutes, and Neutral.
D.2 Hyperparameter search
We fix the number of epochs to 5 and the seed to 1 and perform a hyperparameter search over the learning rates (1e-7, 5e-7, 1e-6, 5e-6, 1e-5, 5e-5) and batch sizes (4, 8, 16), evaluated on the respective validation set. When training on multiple datasets, the overall performance considers each instance equally (instead of averaging dataset scores). The results are shown in Tables 11-14. We train five models using different seeds (1-5) based on the best-performing hyperparameters and report the averaged test performance in Table 15.
Model lr batch-size Acc. F1-score
deberta-v3-large 1e-07 4 49.5 24.7
deberta-v3-large 1e-07 8 49.5 24.7
deberta-v3-large 1e-07 16 49.5 24.7
deberta-v3-large 5e-07 4 50.5 26.2
deberta-v3-large 5e-07 8 49.5 24.7
deberta-v3-large 5e-07 16 49.0 22.1
deberta-v3-large 1e-06 4 58.5 43.7
deberta-v3-large 1e-06 8 53.0 31.6
deberta-v3-large 1e-06 16 50.0 22.2
deberta-v3-large 5e-06 4 90.5 88.6
deberta-v3-large 5e-06 8 88.5 86.4
deberta-v3-large 5e-06 16 82.5 79.6
deberta-v3-large 1e-05 4 91.5 90.4
deberta-v3-large 1e-05 8 91.5 90.2
deberta-v3-large 1e-05 16 88.5 86.8
deberta-v3-large 5e-05 4 50.0 22.2
deberta-v3-large 5e-05 8 50.0 22.2
deberta-v3-large 5e-05 16 82.5 80.7
Table 11: Hyperparameter search reported on our SciFact validation split. We fix the random seed to $1$ and number of epochs to $5$.
Model lr batch-size Acc. F1-score
deberta-v3-large 1e-07 4 69.0 40.8
deberta-v3-large 1e-07 8 69.0 40.8
deberta-v3-large 1e-07 16 69.0 40.8
deberta-v3-large 5e-07 4 75.4 61.8
deberta-v3-large 5e-07 8 69.0 40.8
deberta-v3-large 5e-07 16 69.0 40.8
deberta-v3-large 1e-06 4 85.0 83.2
deberta-v3-large 1e-06 8 85.4 83.6
deberta-v3-large 1e-06 16 69.0 40.8
deberta-v3-large 5e-06 4 90.5 89.1
deberta-v3-large 5e-06 8 90.2 88.8
deberta-v3-large 5e-06 16 89.0 87.4
deberta-v3-large 1e-05 4 88.8 87.4
deberta-v3-large 1e-05 8 90.9 89.3
deberta-v3-large 1e-05 16 88.8 87.1
deberta-v3-large 5e-05 4 69.0 40.8
deberta-v3-large 5e-05 8 69.0 40.8
deberta-v3-large 5e-05 16 89.7 88.3
Table 12: Hyperparameter search reported on the CovidFact validation split. We fix the random seed to $1$ and number of epochs to $5$.
Model lr batch-size Acc. F1-score
deberta-v3-large 1e-07 4 51.6 29.4
deberta-v3-large 1e-07 8 49.8 29.3
deberta-v3-large 1e-07 16 46.7 30.1
deberta-v3-large 5e-07 4 72.3 65.6
deberta-v3-large 5e-07 8 66.2 50.2
deberta-v3-large 5e-07 16 64.8 47.5
deberta-v3-large 1e-06 4 80.1 78.2
deberta-v3-large 1e-06 8 76.5 72.0
deberta-v3-large 1e-06 16 69.1 56.0
deberta-v3-large 5e-06 4 86.6 84.6
deberta-v3-large 5e-06 8 85.4 83.4
deberta-v3-large 5e-06 16 86.3 84.5
deberta-v3-large 1e-05 4 85.6 83.6
deberta-v3-large 1e-05 8 85.7 83.4
deberta-v3-large 1e-05 16 85.8 83.6
deberta-v3-large 5e-05 4 51.8 22.7
deberta-v3-large 5e-05 8 51.8 22.7
deberta-v3-large 5e-05 16 51.8 22.7
Table 13: Hyperparameter search reported on the HealthVer validation split. We fix the random seed to $1$ and number of epochs to $5$.
Model lr batch-size Acc. F1-score
deberta-v3-large 1e-07 4 54.6 33.5
deberta-v3-large 1e-07 8 52.9 28.1
deberta-v3-large 1e-07 16 51.7 27.9
deberta-v3-large 5e-07 4 69.6 62.5
deberta-v3-large 5e-07 8 65.1 46.6
deberta-v3-large 5e-07 16 64.0 46.0
deberta-v3-large 1e-06 4 79.4 78.0
deberta-v3-large 1e-06 8 76.5 73.6
deberta-v3-large 1e-06 16 69.7 59.7
deberta-v3-large 5e-06 4 85.2 83.6
deberta-v3-large 5e-06 8 84.7 83.1
deberta-v3-large 5e-06 16 84.9 83.6
deberta-v3-large 1e-05 4 52.3 22.9
deberta-v3-large 1e-05 8 84.8 83.1
deberta-v3-large 1e-05 16 85.1 83.8
deberta-v3-large 5e-05 4 52.3 22.9
deberta-v3-large 5e-05 8 52.3 22.9
deberta-v3-large 5e-05 16 52.3 22.9
Table 14: Hyperparameter search reported on the SciFact, CovidFact, and HealthVer validation split. We fix the random seed to $1$ and number of epochs to $5$.
Overall Per Label (F1)
Datasets Acc. F1 S N R
SciFact 88.9 87.9 92.0 86.0 85.8
HealthVer 82.1 81.4 80.1 86.6 77.4
CovidFact 90.7 89.5 86.0 93.1
All 84.1 83.3 83.8 86.7 79.5
Table 15: Evaluation of the scientific AFC models on the test set of the respective dataset used for training. The F1-score is macro averaged. The results are averaged over five seeds
Appendix E Details on Finding the Kernel of Truth
E.1 Implementation Details
E.1.1 Baselines
As baselines, we randomly shuffle all passages (random) using five different seeds (1-5) or order passages as they appear within the original publication (ordered). We rank passages via BM25888https://pypi.org/project/rank-bm25/ for ranking based on lexical similarity. For pre-processing, we use SpaCy [^9] https://spacy.io/ for tokenization and converting all tokens to lowercase.
E.1.2 Embedding-based approaches
Model Agg. Sent MRR P@1
BioBERT-ST concat all 0.657 0.462
BioBERT-ST mean 1 0.676 0.500
BioBERT-ST mean 2 0.664 0.500
BioBERT-ST mean 3 0.678 0.500
PubMedBERT-ST concat all 0.764 0.654
PubMedBERT-ST mean 1 0.620 0.423
PubMedBERT-ST mean 2 0.661 0.462
PubMedBERT-ST mean 3 0.664 0.462
SapBERT-ST concat all 0.701 0.538
SapBERT-ST mean 1 0.664 0.500
SapBERT-ST mean 2 0.680 0.462
SapBERT-ST mean 3 0.663 0.462
SBERT concat all 0.663 0.500
SBERT mean 1 0.620 0.423
SBERT mean 2 0.655 0.462
SBERT mean 3 0.602 0.385
SPICED IMS concat all 0.699 0.538
SPICED IMS mean 1 0.598 0.346
SPICED IMS mean 2 0.717 0.577
SPICED IMS mean 3 0.673 0.500
INSTRUCTOR concat all 0.726 0.577
INSTRUCTOR mean 1 0.671 0.500
INSTRUCTOR mean 2 0.763 0.654
INSTRUCTOR mean 3 0.747 0.615
Table 16: Hyper-parameter search on the validation split for sentence-embedding models to select passages $S_{j}^{0}\Rightarrow p_{0}$.
We experiment with various sentence-embedding models (in addition to those reported in Table 1). This includes SBERT101010all-mpnet-base-v2 (Reimers and Gurevych, 2019), and sentence transformers fine-tuned on BioBERT (Lee et al., 2020), PubMedBERT (Gu et al., 2021), and SapBERT (Liu et al., 2021) as provided by Deka et al. (2022). For the INSTRUCTOR, we use the following prompts for the INSTRUCTOR (Su et al., 2022), which follows the official templates:
  • Prompt (Claim): “Represent the Scientific claim for retrieving supporting sentences: ”
  • Prompt (Passage Sentences): “Represent the Scientific sentence for retrieval: ”
For all embedding models, we ranked the passages according to the cosine similarity to the embedded claim. For SPICED-IMS, we use the model [^11] https://github.com/copenlu/scientific-information-change provided by Wright et al. (2022) and rank passages according to their IMS between claim and passage. We treat the scientific AFC models from §D as cross-encoder models to jointly encode each claim-passage pair and re-rank passages based on the predicted probability for the label Supported.
Hyperparameter Tuning
We compare two perspectives on how to represent passages. First, the model embeds the entire passage by concatenating all sentences (denoted as concat). Second, we compute the cosine-similarity (or IMS) between the claim and each sentence of the passage individually and rerank passages based on the mean score of the top $k$ sentences with the highest score (denoted as mean). The intuition is that only fractions of a passage may be relevant for the claim, and focusing on only parts of the passage can benefit the ranking performance. We report the performance on the validation set in Table 16 and select the best hyperparameters based on the P@$1$.
E.1.3 LLMs with PRP
MRR after iteration i
Prompt Section Title P@1 i=1 i=2 i=3 All
Qin et al. (2023) no 0.628 0.750 0.778 0.783 0.783
Qin et al. (2023) (claim) no 0.705 0.806 0.833 0.835 0.837
Qin et al. (2023) (convincing) no 0.731 0.825 0.849 0.848 0.848
Qin et al. (2023) (support) no 0.615 0.739 0.758 0.758 0.760
Qin et al. (2023) yes 0.692 0.798 0.824 0.824 0.825
Qin et al. (2023) (claim) yes 0.731 0.822 0.844 0.841 0.843
Qin et al. (2023) (convincing) yes 0.744 0.831 0.858 0.858 0.858
Qin et al. (2023) (support) yes 0.667 0.769 0.789 0.791 0.793
Table 17: Evaluation of different prompts for PRP with Llama2-70B ( 8bit quantization), with random initialization. Averaged over three seeds.
For PRP, we use the claim and two passages as input to the LLM and prompt it to output which passage should be ranked first. Following Qin et al. (2023), we use these outputs to re-rank passages akin to bubble sort. The bubble sort algorithm ensures that the top $k$ elements are ranked at the top after $k$ iterations. We evaluate four prompts on the validation split in Table 17, starting with the prompt provided by Qin et al. (2023) and three deviations tailored towards our task. We further assess whether passages should include or exclude the title and found that including the title generally improves performance. During prompt selection in Table 17 we randomly shuffle all passages prior to reranking them to avoid the impact of the strong positional bias on the validation split (P@$1$: 0.577). Since our metrics are only sensitive to the top-ranked results, PRP can only be completed over a few iterations to reduce costs without affecting the performance much. Following our experiments on the validation split, we only ran PRP for three iterations on the test set. Prompts are listed in §I.1.
E.2 Exhaustive Results on MissciPlus
Annotated Passages (closed) All Passages (open)
Model P@1 MRR P@1 MRR HasPos@3
Random 0.360 0.566 0.096 0.209 0.213
Ordered 0.480 0.658 0.320 0.443 0.507
BM25 0.547 0.705 0.387 0.539 0.627
SBERT (Reimers and Gurevych, 2019) 0.400 0.631 0.280 0.460 0.547
PubMedBERT ST (Deka et al., 2022) 0.440 0.652 0.240 0.442 0.573
BioBERT ST (Deka et al., 2022) 0.547 0.712 0.427 0.582 0.680
SapBERT ST (Deka et al., 2022) 0.480 0.672 0.333 0.514 0.627
INSTRUCTOR (Su et al., 2022) 0.573 0.738 0.480 0.631 0.733
SPICED-IMS (Wright et al., 2022) 0.587 0.742 0.533 0.664 0.760
DeBERTav3 SciFact (Wadden et al., 2020) 0.603 0.748 0.389 0.535 0.627
DeBERTav3 CovidFact (Saakyan et al., 2021) 0.517 0.691 0.307 0.450 0.507
DeBERTav3 HealthVer (Sarrouti et al., 2021) 0.608 0.765 0.347 0.516 0.629
DeBERTav3 Scienfic AFC (all) 0.608 0.768 0.349 0.514 0.600
Llama2-70B (Touvron et al., 2023) PRP (it=3) 0.711 0.830
Llama3-8B PRP (it=3) 0.729 0.850
GPT3.5 PRP (it=3) 0.671 0.815
GPT4 (OpenAI, 2023) PRP (it=3) 0.742 0.850
Table 18: Ranking performance to find the passages based on which the claim was made ($S_{j}^{0}$) over all 75 test instances where $p_{0}$ was linked to at least one passage.
Table 18 provides a list of all evaluated models, including additional sentence embedding models, tasked to find the passage $S^{0}$ based on which the claim was made. We additionally report HasPositives@$3$ (Shaar et al., 2020) as a more relaxed measure that allows an appropriate passage to be ranked within the top three results.
Appendix F Details on Finding Undermining Passages
We only evaluate finding undermining passages in the open subset. Hence, the results represent a lower bound of the system’s performance. Per our hyper-parameter search (§F.1.1 and §F.1.2), we always provide one randomly sampled passage $S^{0}$ to the model, which is necessary to understand the rationale (and reasoning gaps) behind the claim. If no passage $S^{0}$ exists, we use the paraphrased accurate premise from Missci instead. All experiments are averaged over five seeds (1-5).
F.1 Implementation Details
F.1.1 AFC as Rankers
mAP when $S_{j}^{0}$ included as
Train data Strategy No $S_{j}^{0}$ Claim Evid.
SciFact Support 0.304 0.389 0.193
SciFact Refute 0.258 0.251 0.230
SciFact Both 0.348 0.404 0.208
CovidFact Support 0.236 0.369 0.206
CovidFact Refute 0.161 0.150 0.214
HealthVer Support 0.270 0.357 0.235
HealthVer Refute 0.234 0.269 0.241
HealthVer Both 0.273 0.375 0.245
All AFC Support 0.308 0.410 0.188
All AFC Refute 0.270 0.183 0.192
All AFC Both 0.355 0.420 0.191
Table 19: Evaluation of ranking passages based on the predicted probability of supported, refuted or their sum (both) on the validation split.
A key challenge in retrieving passages that indicate reasoning gaps in a zero-shot setting is the absence of directly comparable tasks. For instance, passages containing content that undermines an argument may convey a supporting stance (e.g., if the claim is confirmed based on a small sample size, as in Hasty Generalization). When re-purposing AFC models to predict whether $S_{j}$ points to a reasoning gap in the argument, we experiment with various strategies that aggregate the stance prediction labels from AFC models. Specifically, we measure the predicted probability mass for the labels Supported, Refuted, or their sum [^12] Except when trained on CovidFact, which only exhibits two labels. (both). For each strategy, we experiment with the role of the passage $S^{0}$ based on which the claim was made:
  • No $S^{0}$: The AFC model sees only $\overline{c}$ and $S_{j}$.
  • Claim: The claim is reformulated as an argument by incorporating $S^{0}$ through the expression “$S^{0}$ Therefore: claim”.
  • Evid.: $S^{0}$ is added at the beginning of the evidence passage ($S_{j}$) that is subject to ranking.
Table 19 shows the validation split results. We select the best-performing strategy according to MAP, which always concatenates a randomly sampled passage $S^{0}$ with the claim.
F.1.2 Dense embedding rankers
MAP
Model no $S_{j}^{0}$ $S_{j}^{0}$
BM25 0.283 0.554
SBERT (concat) 0.328 0.522
SBERT (mean-1) 0.333 0.366
SBERT (mean-2) 0.353 0.354
SBERT (mean-3) 0.308 0.342
INSTRUCTOR (concat) 0.388 0.582
INSTRUCTOR (mean-1) 0.404 0.480
INSTRUCTOR (mean-2) 0.389 0.516
INSTRUCTOR (mean-3) 0.368 0.479
SPICED-IMS (concat) 0.384 0.562
SPICED-IMS (mean-1) 0.379 0.429
SPICED-IMS (mean-2) 0.380 0.408
SPICED-IMS (mean-3) 0.323 0.380
PubMedBERT ST (concat) 0.283 0.490
PubMedBERT ST (mean-1) 0.337 0.529
PubMedBERT ST (mean-2) 0.318 0.513
PubMedBERT ST (mean-3) 0.324 0.490
BioBERT ST (concat) 0.270 0.516
BioBERT ST (mean-1) 0.313 0.531
BioBERT ST (mean-2) 0.311 0.536
BioBERT ST (mean-3) 0.280 0.521
SapBERT ST (concat) 0.323 0.484
SapBERT ST (mean-1) 0.312 0.497
SapBERT ST (mean-2) 0.325 0.500
SapBERT ST (mean-3) 0.323 0.500
Table 20: Ranking performance of embedding-based models on the validation instances to select passages linked to fallacies. We select the best model based on the highest MAP.
Using the same ranking models via embeddings or BM25 as in §E, we assess the impact of prepending a passage $S^{0}$ to the claim on the validation split in Table 20. We select the best configuration according to mAP for the test split. For the INSTRUCTOR, we modify the prompt to encode the claim by looking for “refuting” instead of “supporting” sentences (cf. §E.1.2), which resulted in a higher MAP on the validation set.
F.2 Impact of $S^{0}_{j}$ Passages
Figure 9: Recall of undermining passages per fallacy class over the top $k$ ranked passages when removing all $S^{0}$ passages (based on which the claim was made). Only listing fallacies with $\geq 20$ occurrences.
We artificially remove all passages from the ranked results if the passage communicates the content based on which the claim was made (i.e., $S_{j}^{0}$ passages) in Figure 9. Among these passages only, both rankers exhibit different strengths. INSTRUCTOR prefers passages based on which the Biased Sample Fallacy or the Fallacy of Exclusion can be detected. The AFC ranker prioritizes fallacies based on which Hasty Generalization and False Equivalence can be detected. An intuitive explanation is that both fallacies are similar to the support relationship and, hence, to the predicted probability for the label Supported.
F.3 Exhaustive Results on MissciPlus
Passage-wise Fallacy-wise
Model mAP P@1 P@3 P@10 R@1 R@3 R@10
Random 0.205 0.136 0.130 0.124 0.168 0.252 0.438
Ordered 0.286 0.298 0.210 0.145 0.266 0.374 0.497
BM25 0.496 0.617 0.347 0.212 0.413 0.485 0.602
SBERT (Reimers and Gurevych, 2019) 0.520 0.640 0.381 0.221 0.423 0.503 0.609
INSTRUCTOR (Su et al., 2022) 0.541 0.652 0.409 0.222 0.439 0.529 0.613
SPICED-IMS (Wright et al., 2022) 0.524 0.640 0.388 0.219 0.423 0.516 0.595
PubMedBERT ST (Deka et al., 2022) 0.489 0.588 0.360 0.207 0.387 0.548 0.605
BioBERT ST (Deka et al., 2022) 0.491 0.600 0.377 0.200 0.400 0.495 0.570
SapBERT ST (Deka et al., 2022) 0.504 0.619 0.379 0.211 0.411 0.547 0.615
DeBERTaV3 SciFact (Wadden et al., 2020) 0.360 0.326 0.266 0.172 0.264 0.393 0.518
DeBERTaV3 CovidFact (Saakyan et al., 2021) 0.380 0.457 0.260 0.165 0.326 0.407 0.538
DeBERTaV3 HealthVer (Sarrouti et al., 2021) 0.368 0.410 0.284 0.176 0.321 0.426 0.544
DeBERTaV3 Scientific AFC (all) 0.306 0.338 0.263 0.177 0.267 0.400 0.554
Table 21: Ranking results for finding passages linked to fallacies on the test split, using all passages of the misrepresented publication.
We list all results of detecting undermining passages on the MissciPlus test split in Table 21.
Appendix G Details on Argument Reconstruction
G.1 LLM argument reconstruction
Fallacy Level (premise+class) Arg. Level Count
Prompt Passages R@5 P@5 P@1 F1@5 Arg@1 Pred/Arg
p1-basic per-passage 0.079 0.046 0.133 0.306 0.167 2.6
p2-support per-passage 0.083 0.048 0.078 0.296 0.200 2.7
p3-undermine per-passage 0.079 0.047 0.089 0.302 0.178 2.4
p4-connect per-passage 0.176 0.134 0.222 0.359 0.333 1.4
p5-auto per-passage 0.083 0.066 0.144 0.263 0.178 1.4
p6-auto-connect per-passage 0.111 0.100 0.156 0.272 0.256 1.1
p7-passage per-passage 0.093 0.047 0.078 0.357 0.222 4.2
p8-passage-assumptions per-passage 0.162 0.084 0.122 0.379 0.356 4.2
p9-passage-detective per-passage 0.157 0.146 0.222 0.270 0.344 1.1
p10-passage-evaluate per-passage 0.120 0.060 0.111 0.337 0.289 4.8
p11-passage-evaluate-2 per-passage 0.176 0.087 0.178 0.353 0.367 4.6
p12-passage-detective-2 per-passage 0.218 0.113 0.211 0.360 0.444 3.9
p13-passage-evaluate-3 per-passage 0.167 0.083 0.167 0.359 0.300 4.4
p1-basic all passages 0.060 0.053 0.067 0.294 0.144 2.8
p2-support all passages 0.065 0.058 0.044 0.303 0.156 2.8
p3-undermine all passages 0.046 0.040 0.044 0.280 0.111 2.9
p4-connect all passages 0.093 0.162 0.178 0.239 0.211 1.4
p5-auto all passages 0.056 0.070 0.089 0.247 0.133 2.0
p6-auto-connect all passages 0.046 0.110 0.111 0.163 0.111 1.0
p7-passage all passages 0.042 0.024 0.022 0.333 0.100 4.4
p8-passage-assumptions all passages 0.144 0.088 0.089 0.376 0.289 4.2
p9-passage-detective all passages 0.069 0.153 0.133 0.218 0.167 1.1
p10-passage-evaluate all passages 0.130 0.069 0.167 0.377 0.233 4.7
p11-passage-evaluate-2 all passages 0.144 0.077 0.089 0.371 0.256 5.1
p12-passage-detective-2 all passages 0.171 0.105 0.144 0.379 0.400 3.9
p13-passage-evaluate-3 all passages 0.162 0.094 0.124 0.402 0.344 4.4
Table 22: Argument reconstruction prompt-tuning using Llama3-8B-Instruct on the validation split.
We provide the LLM with all passages $S_{j}$ that were linked to at least one reasoning gap (or fallacy), and one passage $S^{0}$ that communicates the content based on which the claim was made. If multiple candidates for $S^{0}$ exist, we randomly select one. Experiments over Missci use the paraphrased $p_{0}$ instead of a randomly sampled passage $S^{0}$, and the publication context $s_{i}$ instead of the linked passages $S_{j}$. We run each experiment over three different seeds (1-3). For Llama3-8B we use a temperature of 0.3. For a fair assessment given the changed requirements for the LLM, we perform extensive prompt-search over all six prompts evaluated for Missci and an additional new prompts (cf. §I.3). During prompt search, we provide the complete fallacy information consisting of the definition, the logical form, and an example (cf. §A; Tables 7-8) within the instructions. All prompts task LLMs to output a ranked list of verbalized fallacious premises and the applied fallacy class. If possible, fallacies outside our inventory were converted; otherwise, they were removed. The results are listed in Table 22. We additionally report the average number of predicted fallacies per argument. This number excludes fallacies that the LLM hallucinated and are invalid (e.g., Contextomizer, Accident, Conflict of Interest, Fallacy of Conclusion) or that are outside of our fallacy inventory (e.g., Fear Mongering, Non Sequitur, Ad Hominem, False Consensus).
G.2 Llama3 Judge $\phi$
Method Prompt Extra F1 Acc.
Majority 0.372 0.593
NLI-S + LR 0.642 0.695
Zeroshot p1 0-shot 0.570 0.581
Zeroshot p2 0-shot 0.636 0.647
Zeroshot p3 0-shot 0.636 0.641
Zeroshot p4 0-shot 0.626 0.629
Zeroshot p5 0-shot 0.562 0.569
ICL p1 8-shot 0.623 0.625
ICL p2 8-shot 0.635 0.639
ICL p3 8-shot 0.620 0.623
ICL p4 8-shot 0.604 0.607
ICL p5 8-shot 0.612 0.617
ICL p1 16-shot 0.656 0.663
ICL p2 16-shot 0.640 0.645
ICL p3 16-shot 0.645 0.651
ICL p4 16-shot 0.656 0.661
ICL p5 16-shot 0.652 0.657
SFT p4 1-epochs 0.671 0.687
SFT p4 2-epochs 0.690 0.711
SFT p4 3-epochs 0.725 0.749
SFT p4 4-epochs 0.751 0.770
SFT p4 5-epochs 0.788 0.798
SFT p4 6-epochs 0.761 0.776
Table 23: Cross-validation evaluation for implementations of the $\phi^{\mathrm{p+f}}$ judge. All prompt-based approaches use Llama3-8B-Instruct as backend.
We experiment with four different prompts (cf. §I.2) with the Llama3-8B-Instruct model as a binary classifier $\phi$ that determines whether two fallacious premises exhibit fallacious reasoning that bridges the identical gap. We use the human evaluation data provided by Glockner et al. (2024a) for training data. We discard all trivially invalid instances where the generated premise $\hat{\overline{p}}_{i}$ was almost identical to the claim or the paraphrased publication context $s_{i}$ by discarding all premise pairs if
$\mathrm{min}\Bigl{[}\mathrm{lev(\hat{\overline{p}}_{i},\overline{c})},\mathrm{lev(\hat{\overline{p}}_{i},\overline{s}_{i})}\Bigr{]}\leq t$
where $\mathrm{lev}$ is the Levensthein distance and the threshold $t=2$. This yielded a total of 168/240 manually annotated premise pairs. To adapt Llama3-8B-Instruct for the task, we perform (1) zero-shot experiments, (2) in-context learning (Brown et al., 2020) (ICL) experiments and (3) supervised fine-tuning (SFT) using QLoRA (Dettmers et al., 2023) and 8bit quantization (Dettmers et al., 2022). We validate each approach using five-fold cross-validation where folds are separated by the arguments, ensuring the LLM is evaluated with premises from unseen arguments. We set the temperature to zero and evaluate all ICL and SFT experiments over three random seeds ($1$,$2$,$3$) to account for different (ordering) of seen instances. The results are listed in Table 23. As a baseline, we report the performance of always predicting the majority label. We further report a baseline using a univariate logistic regression (LR) on top of the automatic NLI-S (Glockner et al., 2024a), which showed the highest correlation with human judgment in Missci. NLI-S uses the predicted entailment probability of a T5 (Raffel et al., 2020) model fine-tuned by Honovich et al. (2022). Rather than only considering the entailment score given the reference text as a premise and generated text as a hypothesis, NLI-S swaps its roles to avoid penalizing the model if it generates more specific text. Without SFT, ICL with 16 shots and the template p4 reached the best performance measured via the F1-score. We consequentially performed SFT using the same prompt, which led to the overall best model with fine-tuning for $5$ epochs with a linear schedule, a learning rate of $45e-$ 4, a batch-size of $4$ and $\alpha=16$, $r=64$, $\mathrm{dropout}=0.2$ for QLoRA. We use these hyperparameters to fine-tune Llama3-8B-Instruct on the entire data and use it as the backend model for $\phi^{\mathrm{f+p}}$.
G.2.1 Prompting Strategies
Prompts using per-passage prompting follow the prompting scheme of Missci and have a dedicated field for the content based on which the claim was made ($p_{0}$ or $S_{j}^{0}$) and for the publication context necessary to detect the fallacy ($s_{i}$ or $S_{j}$). This prompting technique requires $n$ separate prompts for $n$ passages. Specifically, we create $n-1$ prompts using a randomly sampled passage $S^{0}$ (based on which the claim was made) together with each other passage linked to a fallacy, separately. In MissciPlus, the passage $S^{0}$ itself may be linked to fallacies. Hence, we prompt the model again with only the passage $S^{0}$. When selecting the top $k$ results of multiple prompts for the same argument, combine all results, consisting of the fallacious premise and fallacy class, while keeping their ranking information. We then return the top $k$ results based on their prompt-specific rank. We prefer results with different fallacy classes when multiple fallacies share the same rank to avoid sampling the same fallacious reasoning numerous times. Prompts that concatenate all passages within a single prompt follow the holistic view of arguments. We sort all passages based on the order in which they occur in the scientific publication and only prompt the LLM once per argument.
G.2.2 Performance on Missci with only linked publication context
LLM Info R@5 ($\phi^{\mathrm{f+p}}$) R@5 ($\phi^{\mathrm{f}}$) Arg@$1$ ($\phi^{\mathrm{f+p}}$)
DLE 0.255 0.520 0.516
Llama3-8B DL 0.237 0.453 0.520
DE 0.220 0.470 0.468
LE 0.251 0.492 0.504
DLE 0.217 0.456 0.464
GPT-3.5 DL 0.198 0.438 0.413
DE 0.229 0.461 0.500
LE 0.213 0.417 0.464
DLE 0.327 0.472 0.619
GPT-4 Turbo DL 0.280 0.458 0.560
DE 0.294 0.500 0.571
LE 0.299 0.514 0.583
Table 24: Argument reconstruction performance only over fallacies that could be detected based on linked passages (or based on the accurate premise alone).
We report the argument reconstruction of the evaluated LLMs in Table 24. This evaluation only considers a prediction of an LLM based on the paraphrased information $s_{i}$ if a passage $S_{j}$ that communicates the same information exists, and the same fallacy could have been detected based on $S_{j}$. This serves as a complementary comparison between the performance on Missci and MissciPlus, which removes benefits over Missci due to additional information.
Appendix H AFC applied on MissciPlus
H.1 Stance Predictions per Fallacy Class
Figure 10: Distribution over predicted fact-checking labels per fallacy class, which was linked to the passage used as evidence, when predicting the veracity label. Results are averaged over five seeds for each fine-tuned AFC model.
We aim to understand how passages that point to different reasoning gaps (and hence different fallacy classes) affect the stance prediction of scientific AFC models. Figure 10 visualizes for each AFC model and fallacy class the distribution over the predicted veracity labels of Supported, Refuted and NEI. Regardless of the fallacy class, rarely any passage is identified as refuting the claim. Most models vary between the labels Supported and NEI with only minor clear trends among different fallacy classes. Interestingly, the model trained on CovidFact, which had the best misinformation detection rate according to Table 5 (as it does not know the label NEI) almost always tends to predict Supported more frequently than Refuted over these passages that exhibit fallacious reasoning behind the claim.
H.2 Binary Fallacy Detection by LLMs
Detected a fallacy
LLM MissciPlus True Claims
Llama2-70B 0.988 0.993
Llama3-8B 0.877 0.783
GPT-3.5 0.853 0.923
Table 25: Binary fallacy evaluation over misinformation from MissciPlus and correct claims from HealthVer and CovidFact.
We adjusted our prompt (cf. §I.4) to allow LLMs to select no fallacy if they consider a claim correct and empirically evaluate three LLMs in how well they detect misinformation in MissciPlus or correct information in the 100 selected claims from HealthVer and CovidFact. Since true claims only come with one evidence passage, we also prompt the LLMs once only for misinformation using the concatenated relevant passages (cf. all passages in §6.2.1). Table 25 reports the ratio of claims identified as fallacious (the LLM did not specifically say that “no fallacy” exists and found at least one fallacy) across both datasets averaged over three seeds. We do not discern which fallacy was detected. Using the prompts to reconstruct fallacious arguments, the LLMs tend to find fallacies in all claims regardless of whether they are correct. We note that our claims are designed to correctly verbalize and identify all applied fallacies in MissciPlus. This follows previous works on fallacy detection (Da San Martino et al., 2019; Jin et al., 2022; Alhindi et al., 2022) that focused on which fallacy was applied rather than determining if a fallacy was applied. Future work may explore how to detect fallacies and correct claims better.
Appendix I Prompts
I.1 PRP Prompts
The prompt template for PRP ranking is shown in Figure 11 with the different [INSTRUCTION] listed in Table 26.
[INSTRUCTIONS] Passage A: [PASSAGE-A] Passage B: [PASSAGE-B] Output Passage A or Passage B:
Figure 11: Prompt template for PRP prompting.
Name Instructions
Qin et al. (2023) Given a query [CLAIM], which of the following two passages is more relevant to the query?
Qin et al. (2023) (claim) Given a claim [CLAIM], which of the following two passages is more relevant to the claim?
Qin et al. (2023) (convincing) Given a claim [CLAIM], which of the following two passages constitutes more convincing evidence for the claim?
Qin et al. (2023) (support) Given a claim [CLAIM], is it more likely that someone made the claim based on passage A or based on passage B?
Table 26: Evaluated instructions used for PRP prompting.
I.2 Premise judge $\phi^{\mathrm{p+f}}$
[INSTRUCTIONS] Premises: 1: "[GEN-PREMISE]" 2: "[GOLD-PREMISE]" Question: Do both premises use the same flawed reasoning (fallacy)?
Figure 12: Prompt for the judge $\phi^{\mathrm{p+f}}$.
Name Instructions
p1 You are given two premises that exhibit some reasoning of a larger argument. Both premises apply a fallacy. Your task is to select whether the reasoning on an abstract level is identical. If one premise is more specific, they can still apply the same false reasoning. Provide your answer in the first line of your response. Answer with “match” if both premises apply the same false reasoning. Answer with “no-match” if they apply different false reasoning.
p2 I’ll present you with two premises, each containing a fallacy in their reasoning. Analyze both statements: Your task: Determine if the core flawed logic behind the fallacies in both statements is identical. Respond with “match” if the underlying reasoning is the same, even if the specifics differ. Respond with “no-match” if they represent different fallacies.
p3 Task: Analyze both premises and determine if they commit the same type of fallacy.
Answer: (match / no-match)
p4 Determine whether two premises exhibit identical false reasoning, regardless of their specificity. Provide your answer in the first line of your response. Answer with “match” if both premises apply the same false reasoning. Answer with “no-match” if they apply different false reasoning.
p5 Task: Analyze both premises and determine if they apply a similar reasoning regardless of specificity.
Answer: (match / no-match)
Table 27: Instructions used for the judge $\phi^{\mathrm{p+f}}$.
The prompt template for the judge $\phi^{\mathrm{p+f}}$ with ICL and SFT approaches is shown in Figure 12. The instructions are listed in Table 27.
I.3 Argument reconstruction prompts
[FALLACY INFORMATION] Task: Examine the following fallacious argument: Passage 1: "[PASSAGE S0]" Passage 2: "[PASSAGE Sj]" Claim: "[CLAIM]" Passages 1 and 2 are sourced from the same credible scientific document. The claim is based on the content of both passages. Your task is to identify and verbalize the fallacious reasoning as the fallacious premise necessary to support the claim given the content of passages 1 and 2. This reasoning should effectively support the claim, ensuring that the passages do not undermine the claim as a valid conclusion. Only consider fallacies from the provided fallacy inventory. Present each fallacious premise along with the applied fallacy class in this format: Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>. If multiple applicable fallacies exist, list all fallacious premises with the applied fallacy classes in order of relevance (most to least relevant).
Figure 13: Prompt template for argument reconstruction p7-passage.
[FALLACY INFORMATION] Task: Examine the following fallacious argument: Passage 1: "[PASSAGE S0]" Passage 2: "[PASSAGE Sj]" Claim: "[CLAIM]" Passages 1 and 2 are sourced from the same credible scientific document. The claim is based on the content of both passages. Your task is to identify and verbalize the fallacious reasoning (the hidden assumptions) as the fallacious premise necessary to support the claim given the content of passages 1 and 2. This reasoning should effectively support the claim, ensuring that the passages do not undermine the claim as a valid conclusion. Only consider fallacies from the provided fallacy inventory. Present each fallacious premise along with the applied fallacy class in this format: Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>. If multiple applicable fallacies exist, list all fallacious premises with the applied fallacy classes in order of relevance (most to least relevant).
Figure 14: Prompt template for argument reconstruction p8-passage-assumptions.
[FALLACY INFORMATION] Challenge: You’ve been assigned a detective mission in the world of scientific arguments! The Case: An argument has been made based on two passages from a credible scientific document. However, there’s a hidden flaw in the reasoning. Your job is to uncover this hidden assumption – the "fallacious premise" – that makes the argument illogical. The Evidence: Passage 1 (This is the first piece of information from the scientific document.): "[PASSAGE S0]" Passage 2 (This provides additional details from the same document.): "[PASSAGE Sj]" Claim (This is the conclusion drawn from both passages): "[CLAIM]" The Tools: Fallacy Inventory: You have access to a list of common fallacies (errors in reasoning). Only use these fallacies! Your Mission: 1. Analyze the passages and the claim. 2. Identify the hidden assumption that’s needed for the claim to follow logically from the passages. 3. Formulate this hidden assumption as a clear "fallacious premise." 4. Identify the specific type of fallacy from the fallacy inventory that best explains this hidden assumption. Deliverables: Present your findings in this format: "Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>." If multiple applicable fallacies exist, list all fallacious premises with the applied fallacy classes in order of relevance (most to least relevant). Only consider fallacies from the provided fallacy inventory. Remember: The passages come from a credible scientific document, so the information itself is likely true. The fallacy lies in how the information is used to support the claim. Focus on the hidden assumption(s) needed to bridge the gap between the passages and the claim.
Figure 15: Prompt template for argument reconstruction p9-passage-detective.
[FALLACY INFORMATION] Task:
This activity focuses on identifying weaknesses in scientific arguments based on source materials.
Materials:
Passage 1: "[PASSAGE S0]"
Passage 2: "[PASSAGE Sj]" Claim: "[CLAIM]" The claim is derived from the content of both passages. Both passages stem from the same credible scientific publication. Objective:
Analyze the relationship between the provided passages and the claim. Identify any assumptions or gaps in reasoning that might weaken the argument’s validity.
Instructions:
1. Read Passage 1, Passage 2, and the claim carefully.
2. Consider how the information in the passages connects to the claim. 3. Identify any missing information or hidden assumptions that would be necessary for the claim to logically follow from the passages. 4. Formulate these missing pieces as clear "fallacious premises." 5. Using the provided fallacy list, identify the type of fallacy associated with each fallacious premise. Output:
Present your findings in this format:
"Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>." If multiple applicable fallacies exist, list them in order of relevance (most to least relevant). Only consider fallacies from the provided fallacy inventory. Important Note:
The passages are sourced from a credible scientific document, so the information itself is most likely accurate. The focus is on how the information is used to support the claim, not questioning the scientific content.
Figure 16: Prompt template for argument reconstruction p10-passage-evaluate.
[FALLACY INFORMATION] Task:
This activity focuses on identifying weaknesses in scientific arguments based on source materials.
Materials:
Passage 1: "[PASSAGE S0]"
Passage 2: "[PASSAGE Sj]" Claim: "[CLAIM]" The claim is derived from the content of both passages. Both passages stem from the same credible scientific publication. Objective:
Analyze the relationship between the provided passages and the claim. Identify any assumptions or gaps in reasoning that might weaken the argument’s validity.
Instructions:
1. Read Passage 1, Passage 2, and the claim carefully.
2. Consider how the information in the passages connects to the claim. 3. Identify any missing information or hidden assumptions that would be necessary for the claim to logically follow from the passages. 4. Formulate these missing pieces as clear "fallacious premises." 5. Using the provided fallacy list, identify the type of fallacy associated with each fallacious premise. Output:
Present your findings in this format: "Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>." If multiple applicable fallacies exist, list all fallacious premises with the applied fallacy classes in order of relevance (most to least relevant). Only consider fallacies from the provided fallacy inventory.
Important Note:
The passages are sourced from a credible scientific document, so the information itself is most likely accurate. The focus is on how the information is used to support the claim, not questioning the scientific content.
Figure 17: Prompt template for argument reconstruction p11-passage-evaluate-2.
[FALLACY INFORMATION] Challenge: You’ve been assigned a detective mission in the world of scientific arguments! The Case:
An argument has been made based on two passages from a credible scientific document. However, there’s a hidden flaw in the reasoning. Your job is to uncover this hidden assumption – the "fallacious premise" – that makes the argument illogical.
The Evidence:
Passage 1 (This is the first piece of information from the scientific document.): "[PASSAGE S0]"
Passage 2 (This provides additional details from the same document.): "[PASSAGE Sj]" Claim (This is the conclusion drawn from both passages): "[CLAIM]" The Tools:
Fallacy Inventory: You have access to a list of common fallacies (errors in reasoning). Only use these fallacies!
Your Mission: 1. Analyze the passages and the claim. 2. Identify the hidden assumption that’s needed for the claim to follow logically from the passages. 3. FOR EACH identified flaw: a. Formulate this hidden assumption as a clear "fallacious premise." b. Using the provided fallacy inventory, identify the specific fallacy class from the provided fallacy inventory that best explains this fallacious premise. c. Output each identified flaw in a separate line following this format: "Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>." Deliverables:
Present your findings in this format:
"Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>." If multiple applicable fallacies exist, list all fallacious premises with the applied fallacy classes in order of relevance (most to least relevant). Only consider fallacies from the provided fallacy inventory. Remember:
The passages come from a credible scientific document, so the information itself is likely true. The fallacy lies in how the information is used to support the claim. Focus on the hidden assumption(s) needed to bridge the gap between the passages and the claim.
Figure 18: Prompt template for argument reconstruction p12-passage-detective-2.
[FALLACY INFORMATION] Task:
This activity focuses on identifying weaknesses in scientific arguments based on source materials.
Materials:
Passage 1: "[PASSAGE S0]"
Passage 2: "[PASSAGE Sj]" Claim: "[CLAIM]" The claim is derived from the content of both passages. Both passages stem from the same credible scientific publication. Objective:
Analyze the relationship between the provided passages and the claim. Identify any assumptions or gaps in reasoning that might weaken the argument’s validity.
Your Mission:
1. Analyze the passages and the claim.
2. Identify the hidden assumption that’s needed for the claim to follow logically from the passages. 3. FOR EACH identified flaw: a. Formulate this hidden assumption as a clear "fallacious premise." b. Using the provided fallacy inventory, identify the specific fallacy class from the provided fallacy inventory that best explains this fallacious premise. c. Output each identified flaw in a separate line following this format: "Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>." Output:
Present your findings in this format:
"Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>." If multiple applicable fallacies exist, list all fallacious premises with the applied fallacy classes in order of relevance (most to least relevant). Only consider fallacies from the provided fallacy inventory. Important Note:
The passages are sourced from a credible scientific document, so the information itself is most likely accurate. The focus is on how the information is used to support the claim, not questioning the scientific content.
Figure 19: Prompt template for argument reconstruction p13-passage-detective-3.
We take the p1-p6 prompts from Missci as-is. For all-passages prompting we concatenate all passages (except for one $S^{0}_{j}$ passage), treating them as publication context in the Missci prompt, and treating the left-out $S^{0}_{j}$ as the accurate premise. The per-passage prompt templates for p7-p13 are shown in Figures 13-19., where we always replace [PASSAGE S0] with a randomly sampled passage $S_{j}^{0}$ and [PASSAGE Sj] with a passage $S_{j}$ linked to a fallacy. For all, we present the fallacy information consisting of the fallacy definition, logical form and example from literature identically to Missci prompts. The all-passage prompts are minimally edited from these prompts and have one dedicated field for all selected passages. All used prompts are provided within our repository.
I.4 LLM as AFC prompts
Write a fact-checking article about the claim that "[CLAIM]" Conclude your fact-checking article with a verdict in a single line: "Verdict: (your verdict label)". Choose one of the following verdicts: True, Unknown, False.
Figure 20: Prompt to generate a fact-checking article given no evidence.
Claim: "[CLAIM]" To the best of your knowledge, what is the veracity of the claim? Provide a thorough explanation supporting your decision, select one of the answers (True, False, Unknown) and output the veracity of the claim in a single line: "Veracity: (your veracity label)".
Figure 21: Prompt to ask for the veracity of the claim given no evidence.
Study: [EVIDENCE] Rate if the claim "[CLAIM]" is correct based on the study. Answer with "Verdict: (Correct / Incorrect / Not Enough Information)"
Figure 22: Prompt to ask for the veracity of the claim given evidence.
[FALLACY INFORMATION] Challenge:
You’ve been assigned a detective mission in the world of scientific arguments!
The Case:
An argument has been made based on two passages from a credible scientific document. Your job is to assess the logical reasoning behind this argument.
The Evidence:
Passages (The used information from the scientific document): "[EVIDENCE]" Claim (This is the conclusion drawn from the passages): "[CLAIM]"
The Tools:
Fallacy Inventory: You have access to a list of common fallacies (errors in reasoning). Only use these fallacies!
Your Mission:
1. Analyze the passages and the claim.
2. Evaluate whether the claim logically follows from the information presented in the passages. 3. IF you identify flaws in the reasoning, pinpoint all hidden assumptions that’s needed for the claim to follow logically. 4. FOR EACH identified flaw: a. Formulate this hidden assumption as a clear "fallacious premise." b. Using the provided fallacy inventory, identify the specific fallacy class from the provided fallacy inventory that best explains this fallacious premise. c. Output each identified flaw in a separate line following this format: "Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>." 5. IF the claim logically follows from the passages (no fallacy present), then simply output: "No Fallacy" Deliverables:
If the argument is sound (if no fallacy or only minor fallacies exist), clearly state "No Fallacy" in the output.
If a fallacy is present, present your findings in this format: "Fallacious Premise: <fallacious premise>; Applied Fallacy Class: <applied fallacy class>." If multiple applicable fallacies exist, list all fallacious premises with the applied fallacy classes in order of relevance (most to least relevant). Only consider fallacies from the provided fallacy inventory. Remember:
The passages come from a credible scientific document, so the information itself is likely true. The fallacy lies in how the information is used to support the claim. Focus on the hidden assumption(s) needed to bridge the gap between the passages and the claim. If the scientific document supports the claim, no critical fallacy is applied.
Figure 23: Fallacy generation prompt that can predict “no fallacy”.
The prompts used to directly predict the veracity of claims using LLMs are shown in Figures 20-22. We replace [EVIDENCE] with the concatenated evidence passages. The adapted fallacy generation prompt which allows to output that no fallacy exists, is shown in Figure 23.
Appendix J Reproducibility
All experiments with Llama2 or Llama3 used the instruction-tuned LLM and were performed on 80GB A100 GPUs. We always used the 70B and 8B models for Llama2 and Llama3, respectively. All experiments with LLMs are averaged over three seeds. The only exception is the experiments with GPT-4 Turbo in Table 3, which we only ran once due to computational costs. We used the API version 2023-10-01-preview for GPT-3.5 (model version: 0613) [^13] We used gpt-35-turbo-16k in the all passages prompting experiments and gpt-35-turbo in all other experiments, GPT-4 (model version: 0613) and GPT-4 Turbo (model version: 1106-Preview)
We used Grammarly [^14] https://app.grammarly.com/ in writing this paper.